Determine whether the lines through each pair of points are parallel. Thell how you determined the answer.
(0,1) and (3,7)
(1,1) and(2,3)

I got that they were parallel

Answer by Mathtut(3670)

slope=7-1/3-0=6/3=2---
:
slope=3-1/2-1=2
:
slopes are equal and different y intercepts... therefore they are parallel